# Relaxation of first-class constraints and the quantization of gauge theories: from “matter without matter” to the reappearance of time in quantum gravity

### v1

@@ Abstract
We make a conceptual overview of a particular approach to the initial-value problem in canonical gauge theories. We stress how the first-class phase-space constraints may be relaxed if we interpret them as fixing the values of new degrees of freedom. This idea goes back to Fock and Stueckelberg, leading to restrictions of the gauge symmetry of a theory, and it corresponds, in certain cases, to promoting constants of Nature to physical fields. Recently, different versions of this formulation have gained considerable attention in the literature, with several independent iterations, particularly in classical and quantum descriptions of gravity, cosmology, and electromagnetism. In particular, in the case of canonical quantum gravity, the Fock–Stueckelberg approach is relevant to the so-called problem of time. Our overview recalls the work of Fock and Stueckelberg and its physical interpretation with the aim of conceptually unifying the different iterations of the idea that appear in the literature and of motivating further research.
@@ I Introduction
Fock and, more explicitly, Stueckelberg proposed a relaxation of the relativistic particle’s mass-shell constraint, which is the particle’s version of Eq. (1) (the mass-shell constraint follows from diffeomorphism invariance on the worldline, also referred to as time reparametrization invariance Chataig:Thesis ). The lack of a mass-shell constraint simply implies that, instead of fixing the particle’s (rest) mass to be a certain constant (and thereby restricting the particle’s motion to a timelike worldline), its mass should rather be determined by the initial values of the particle’s four-momentum, so that different trajectories could yield different masses. Moreover, the trajectories need no longer be timelike, a phenomenon that Stueckelberg used to describe pair creation and annhiliation by identifying antiparticles with particles going backwards in Minkowski time.
@@
Unimodular gravity theories have attracted a considerable amount of attention in recent years. As is well-known, an approach to unimodular gravity can be obtained by the requirement that the theory be invariant with respect to a restricted class of diffeomorphisms, the so-called volume-preserving, area-preserving or divergenceless diffeomorphisms Arnold ; Wolski ; Bose ; Arakelian ; Floratos ; Bars ; Wit ; closed ; Kugo ; i.e., by the diffeomorphisms generated by the divergenceless transverse vector fields. In such a formalism, the Hamiltonian constraint (1) is not satisfied at all points, but its value is related to an integration constant that plays the role of the cosmological constant. As in the case of the Fock–Stueckelberg formalism for the relativistic particle, the approaches to unimodular gravity can be seen as particular modifications of general relativity, whereby one adds new degrees of freedom and modifies the local symmetry group, which leads to a relaxation of the original constraints.
Evidently, this type of modification is not restricted to gravitational theories. In principle, one can apply the Fock–Stueckelberg reasoning to any first-class initial-value constraint, i.e., to any canonical gauge theory. We will variously refer to such treatments of canonical gauge theories as the Fock–Stueckelberg theory, approach, mechanism, or method. One simple example can be done for electromagnetism, as discussed by the authors of Kaplan and Kaplan1 in the article Kaplan2 , where they relaxed the standard vacuum Gauss-law constraint, ${\rm div}\vec{E}=0$, replacing it by
@@
Although the formalism proposed by the authors of Kaplan ; Kaplan1 ; Kaplan2 may certainly lead to a well-defined quantum theory of interest, it is important to emphasize that the relaxation of constraints is by no means a mandatory step in the quantization of gauge theories. This is evidenced by the success of quantum electrodynamics, not to mention the standard model of particle physics, and by the powerful formalism of quantization à la Batalin–Vilkovisky (BV), Batalin–Fradkin–Vilkovisky (BFV), and Becchi–Rouet–Stora–Tyutin (BRST) (see, e.g., HT:book ). In this way, although it may be easier to quantize a theory without (some of the) constraints, the relaxation of constraints is not a prerequisite for a well-defined quantum theory (at least given our current knowledge of quantum foundations). Rather, as was already mentioned, the relaxation of constraints defines a different classical theory. Both theories, constrained and unconstrained, may be quantized and each may lead to different phenomenology due to the presence or absence of new degrees of freedom (e.g., presence or absence of a mass field in the traditional and Fock–Stueckelberg treatments of the relativistic particle).
@@
Furthermore, it is true that there is a distinction between Yang–Mills symmetries and diffeomorphism invariance but this need not imply that the standard treatment of the constraints in general relativity will lead to a pathological quantum theory. The Yang–Mills symmetries are internal, as they do not concern a change of spacetime point (the argument of fields), whereas diffeomorphisms are external as they “move points.” However, by taking an active view of diffeomorphisms, in which the original and transformed fields may differ in functional form but not in argument, this distinction seems less crucial. In fact, both types of local symmetries can be treated in formally the same way when analyzing the invariance of the action functional, provided one adds suitable boundary terms Vergara . It thus seems premature to a priori dismiss the standard quantization of general relativity’s constraints. The task of finding such a non-perturbative (and possibly ultraviolet complete) theory is surely daunting but not necessarily inconsistent. In this way, the Fock–Stueckelberg approach is not inevitable in quantum gravity and cosmology.
@@
Evidently, however, the necessity of going beyond the standard models of particle physics and cosmology may motivate us to consider the elegant method of Fock and Stueckelberg as a means of accounting for new physics. Although this was indeed the spirit of the articles Kaplan ; Kaplan1 ; Kaplan2 , and of the many other preceding articles cited above, an overview with an emphasis on the physical meaning of the Fock–Stueckelberg theory is long overdue, due to the frequent appearance of versions of this approach in the literature. In the present article, we propose such an overview.
@@
The article is divided into five sections. In Sec. II, we quickly review the fundamentals of canonical gauge theories, in order to subsequently present the Fock–Stueckelberg approach to general gauge theories in Sec. III. This section is then subdivided into several particular cases and examples of interest. In Sec. III.2, we discuss how promoting constants of Nature to spacetime fields can be seen as an equivalent approach to the Fock–Stueckelberg theory under certain conditions. The constraint relaxation for “partially Abelianized” systems is then presented in Sec. III.3, and subsequently we turn our attention to mechanical examples in Secs. III.4 and III.5, where we examine the case of time reparametrization invariance and relativistic particles. The Fock–Stueckelberg approach to electromagnetism is presented in Sec. III.6, whereas the application to general relativity is featured in Sec. III.7. Section IV discusses the implications of the Fock–Stueckelberg method to the problem of time in quantum gravity, and we finish with our conclusions in Sec. V.
@@ II Canonical gauge theories in a nutshell
We begin with a brief review of canonical gauge theories before we discuss a general presentation of what we call the Fock–Stueckelberg approach. Their original work on the relativistic particle will be discussed as a particular example of the general framework.
@@ II.1 Constraint algebra and gauge transformations
Let us consider a set of fields $z^{\mu}(\tau,\mathbf{x})$ ($\mu=1,\ldots,Z$) in $d+1$ spacetime dimensions. We omit the spatial dependence by adopting DeWitt’s compact notation DeWittFields in which the indices also include a spatial dependence, $z^{\mu}(\tau,\mathbf{x})\equiv z^{\mu}(\tau)\equiv z^{\mu}$. [^1] We use the sign $\equiv$ to denote equivalent notations. It is also used when two functions are identically equal (i.e., equal at every point). We will use the sign $\coloneqq$ to denote the definition of a quantity. Repeated indices are taken to imply a tacit summation over their discrete values together with a spatial integration. In the case $d=0$ (mechanics), there is of course no spatial dependence and no spatial integrations. We assume that the fields $z^{\mu}$ are bosonic for simplicity. The field-space spanned by $z^{\mu}$ will be referred to as the phase space.
@@
$S[z,\lambda]=\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\left\{\mathcal{A}_{\mu}(z)\dot{z}^{\mu}-H_{0}(\tau;z]-\lambda^{a}\phi_{a}(\tau;z)\right\}\,,$ (4)
where a dot indicates a derivative with respect to time $\tau$, $\mathcal{A}_{\mu}$ is the symplectic potential, and $\lambda^{a}$ are Lagrange multipliers ($a=1,\ldots,N\leq Z$). The field equations of $z^{\mu}$ are [^3] The derivatives with respect to $z^{\mu}$ are taken in the functional sense if $d>0$. We recall that, for a local functional $\mathcal{F}$ of the field $z(\tau,\mathbf{x})$, which is written as $\mathcal{F}(\tau,\mathbf{x};z]=\int\mathrm{d}^{3}yF\left(\tau,\mathbf{x},\mathbf{y};z(\tau,\mathbf{y}),\mathrm{D}z(\tau,\mathbf{y}),\ldots,\mathrm{D}^{n}z(\tau,\mathbf{y})\right)$ with $F$ being a scalar density in $\mathbf{y}$ and $\mathrm{D}^{n}$ being the $n$-th spatial covariant derivative, the functional derivative $\delta\mathcal{F}/\delta z(\tau,\mathbf{x})$ of $\mathcal{F}$ relative to a variation $z(\tau,\mathbf{x})\mapsto z(\tau,\mathbf{x})+\varepsilon\eta(\tau,\mathbf{x})$ is obtained from $\delta\mathcal{F}=\varepsilon\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\mathcal{F}(\tau,\mathbf{x};z+\varepsilon\eta]=:\varepsilon\int\mathrm{d}^{3}x\ \frac{\delta\mathcal{F}}{\delta z(\tau,\mathbf{x})}\eta(\tau,\mathbf{x})$. This leads to $\delta z(\tau,\mathbf{x})=\varepsilon\eta(\tau,\mathbf{x})$, as it should.
@@
are the symplectic form and the total Hamiltonian, respectively. We assume that $\Omega_{\mu\nu}$ is invertible, so that $Z$ must be even (otherwise $\Omega_{\mu\nu}$ would be singular) and $\Omega^{\mu\nu}\Omega_{\nu\rho}=\delta^{\mu}_{\rho}$. [^4] If $d>0$, the delta symbol corresponds to a multiplication of a Kronecker delta for the discrete variables and a Dirac delta distribution for the continuous spatial coordinates. Then, Eq. (5) can be rewritten as
@@
In this way, the field equations associated with $z^{\mu}$ can be written compactly as $\dot{z}^{\mu}=\{z,H\}$. On the other hand, the field equations associated with the multipliers $\lambda^{a}$ are constraints on the fields:
$\phi_{a}(\tau;z)=0\ .$ (9)
@@
We can now make a series of simplifying assumptions that will be sufficient for our purposes. We assume that the constraint functions do not depend explicitly on time,
$\frac{\partial\phi_{a}}{\partial\tau}=0\ ,$ (10)
@@
and that the $\phi_{a}$ functions are all functionally independent and not identically zero, so there is no constraint that can be reduced to another (e.g., $\phi_{2}=\phi_{1}^{2}$). The constraint equations (9) are also assumed to define a hypersurface $\Sigma$ on the space of the $z$ fields. Furthermore, in open regions covering $\Sigma$, the derivatives $\partial\phi_{a}/\partial z^{\mu}|_{\Sigma}$ must not diverge and must define a matrix with a rank equal to the number $N$ of constraints. This regularity condition HT:book implies that the constraint functions can be regarded as coordinates in a neighborhood of $\Sigma$; i.e., we can define local phase-space coordinates $z^{\prime}(z)=(\tilde{z}(z),\phi(z))$. Then, if $f^{\prime}(z^{\prime})\coloneqq f(z(z^{\prime}))$ is such that $f^{\prime}(\tilde{z},\phi=0)=0$, we obtain [^5] See Theorem 1.1 and Appendix 1.A in HT:book .
$f^{\prime}(\tilde{z},\phi)$ $=\int_{0}^{1}\mathrm{d}\ell\ \frac{\partial}{\partial\ell}f^{\prime}(\tilde{z},\ell\phi)$ (11)
$=\phi_{a}\int_{0}^{1}\mathrm{d}\ell\ \frac{1}{\ell}\frac{\partial}{\partial\phi_{a}}f^{\prime}(\tilde{z},\ell\phi)$
$=:v^{a}(\tilde{z},\phi)\phi_{a}\ .$
Thus, if a phase-space function $f$ vanishes on $\Sigma$, it can be written as $f(z)=v^{a}(z^{\prime}(z))\phi_{a}(z)$ for some functions $v^{a}$, which are locally defined by Eq. (11).
@@
We use Dirac’s ‘weak equality’ sign $\approx$ to denote identities that hold on $\Sigma$ Dirac2 . [^6] Identities that hold when (all) the field equations are satisfied can be called ‘on-shell identities’. Since every solution to the field equations must satisfy Eqs. (9), then on-shell identities are also weak equalities. In a slight abuse of terminology, we use both weak equalities and on-shell identities as synonyms. In particular, $\phi_{a}\approx 0$. The constraint equations (9) must hold at every instant of time, so that $\dot{\phi}_{a}\approx 0$, otherwise the theory is not consistent. The requirement that $\dot{\phi}_{a}\approx 0$ may lead to new constraints $C_{a}$ that are functionally independent from $\phi_{a}$. In the standard Rosenfeld–Bergmann–Dirac analysis Rosenfeld ; SalisburyRosenfeld ; Dirac3 ; RBD ; HT:book ; Bergmann , the $\phi_{a}$ are called ‘primary’ constraints, as they follow from the structure of the action (4), whereas $C_{a}$ are called ‘secondary’ as they follow from the application of the field equations. [^7] Repeated application of the field equations might lead to ‘tertiary’ constraints and so forth, but all these constraints can be collectively called secondary. For our purposes, it is sufficient to consider that only the primaries $\phi_{a}$ are present; i.e., the conditions $\dot{\phi}_{a}\approx 0$ are satisfied on $\Sigma$ without leading to new (secondary) constraints. In this case, if we demand that $\dot{\phi}_{a}\approx 0$ is satisfied for every arbitrary value of the multipliers $\lambda^{a}$, then $\phi_{a}$ and $H_{0}$ must obey a first-class algebra (one that Poisson-commutes on $\Sigma$). Due to Eq. (11), this implies
@@
Is any such transformation allowed? No. If no preferred frame is to be selected, then the field equations must have the same functional form in all frames. This means that $z^{\mu}\to z^{\mu}+\delta z^{\mu}$, $\lambda^{a}\to\lambda^{a}+\delta\lambda^{a}$ should be a symmetry. Then, Noether’s theorem implies that the Noether charge $G$ must satisfy Pons
@@
If $\mathcal{I}$ is gauge invariant, it does not change under arbitrary gauge transformations, i.e., $\delta\mathcal{I}\approx 0$ for any choice of $\varepsilon^{a}$. Using the independence of $\varepsilon^{a}$ and its derivatives implied by its arbitrariness, we see that Eq. (17) together with $\delta\mathcal{I}\approx 0$ implies the gauge invariance conditions:
@@
A gauge-invariant quantity can be seen as one that is the same in every generalized reference frame. In particular, the constraint functions $\phi_{a}(z)$ are gauge invariant, $\delta\phi_{a}\approx 0$, and thus their functional form does not change under a gauge transformation.

### v2

@@ Abstract
We make a conceptual overview of a particular approach to the initial-value problem in canonical gauge theories. We stress how the first-class phase-space constraints may be relaxed if we interpret them as fixing the values of new degrees of freedom. This idea goes back to Fock and Stueckelberg, leading to restrictions of the gauge symmetry of a theory, and it corresponds, in certain cases, to promoting constants of Nature to physical fields. Recently, different versions of this formulation have gained considerable attention in the literature, with several independent iterations, particularly in classical and quantum descriptions of gravity, cosmology, and electromagnetism. In particular, in the case of canonical quantum gravity, the Fock–Stueckelberg approach is relevant to the so-called problem of time. Our overview recalls and generalizes the work of Fock and Stueckelberg and its physical interpretation with the aim of conceptually unifying the different iterations of the idea that appear in the literature and of motivating further research.
@@ I Introduction
Fock and, more explicitly, Stueckelberg proposed a relaxation of the relativistic particle’s mass-shell constraint, which is the particle’s version of Eq. (1) (the mass-shell constraint follows from diffeomorphism invariance on the worldline, also referred to as time reparametrization invariance Chataig:Thesis ). The lack of a mass-shell constraint simply implies that, instead of fixing the particle’s (rest) mass to be a certain constant (and thereby restricting the particle’s motion to a timelike worldline), its mass should rather be determined by the initial values of the particle’s four-momentum, so that different trajectories could yield different masses. Moreover, the trajectories need no longer be timelike, a phenomenon that Stueckelberg used to describe pair creation and annihilation by identifying antiparticles with particles going backwards in Minkowski time.
@@
Unimodular gravity theories have attracted a considerable amount of attention in recent years. As is well-known, an approach to unimodular gravity can be obtained by the requirement that the theory be invariant with respect to a restricted class of diffeomorphisms, the so-called volume-preserving, area-preserving or divergenceless diffeomorphisms Arnold ; Wolski ; Bose ; Arakelian ; Floratos ; Bars ; Wit ; closed ; Kugo ; i.e., by the diffeomorphisms generated by the divergenceless transverse vector fields. In such a formalism, the Hamiltonian constraint (1) is not satisfied at all points, but the value of the Hamiltonian [the left-hand side of the constraint equation (1)] is related to an integration constant that plays the role of the cosmological constant. As in the case of the Fock–Stueckelberg formalism for the relativistic particle, the approaches to unimodular gravity can be seen as particular modifications of general relativity, whereby one adds new degrees of freedom and modifies the local symmetry group, which leads to a relaxation of the original constraints.
Evidently, this type of modification is not restricted to gravitational theories. In principle, one can apply the Fock–Stueckelberg reasoning to any canonical gauge theory. We will variously refer to such treatments of canonical gauge theories as the Fock–Stueckelberg theory, approach, mechanism, or method. One simple example can be done for electromagnetism, as discussed by the authors of Kaplan and Kaplan1 in the article Kaplan2 , where they relaxed the standard vacuum Gauss-law constraint, ${\rm div}\vec{E}=0$, replacing it by
@@
Although the formalism proposed by the authors of Kaplan ; Kaplan1 ; Kaplan2 may certainly lead to a well-defined quantum theory of interest, it is important to emphasize that the relaxation of constraints is by no means a mandatory step in the quantization of gauge theories. This is evidenced by the success of quantum electrodynamics, not to mention the standard model of particle physics, and by the powerful formalism of quantization à la Batalin–Vilkovisky (BV), Batalin–Fradkin–Vilkovisky (BFV), and Becchi–Rouet–Stora–Tyutin (BRST) (see, e.g., HT:book ). In this way, although it may be easier to quantize a theory without (some of the) constraints, the relaxation of constraints is not a prerequisite for a well-defined quantum theory (at least given our current knowledge of quantum foundations). Rather, as was already mentioned, the relaxation of constraints defines a different classical theory. Both theories, constrained and unconstrained, may be quantized and each may lead to different phenomenology due to the presence or absence of new degrees of freedom (e.g., presence or absence of a mass field in the traditional and Fock–Stueckelberg treatments of the relativistic particle). This is a central point of our analysis.
@@
Furthermore, it is true that there is a distinction between Yang–Mills symmetries and diffeomorphism invariance but this need not imply that the standard treatment of the constraints in general relativity will lead to a pathological quantum theory. The Yang–Mills symmetries are internal, as they do not concern a change of spacetime point (the argument of fields), whereas diffeomorphisms are external as they “move points.” However, by taking an active view of diffeomorphisms, in which the original and transformed fields may differ in functional form but not in argument, this distinction seems less crucial. In fact, both types of local symmetries can be treated in formally the same way when analyzing the invariance of the action functional, provided one adds suitable boundary terms Vergara . It thus seems premature to a priori dismiss the standard quantization of general relativity’s constraints. The task of finding such a (possibly ultraviolet complete) quantization that is non-perturbative is surely daunting but not necessarily impossible. In this way, the Fock–Stueckelberg approach is not inevitable in quantum gravity and cosmology.
@@
Evidently, however, the necessity of going beyond the standard models of particle physics and cosmology may motivate us to consider the elegant method of Fock and Stueckelberg as a means of accounting for new physics. Although this was indeed the spirit of the articles Kaplan ; Kaplan1 ; Kaplan2 , and of the many other preceding articles cited above, an overview with an emphasis on the physical meaning of the Fock–Stueckelberg theory is long overdue, due to the frequent appearance of versions of this approach in the literature. In the present article, we propose such an overview, which not only recalls the original works of Fock and Stueckelberg, but also offers a generalization that encompasses different approaches in the literature.
The title of the present article mentions the relaxation of first-class constraints and the quantization of gauge theories. What is their relation? As discussed above, and as will be argued throughout, this relaxation is not needed to achieve a consistent and general quantization of gauge theories, despite what some of the literature may imply. We will show below that the relaxation of constraints à la Fock and Stueckelberg is a mechanism to define new classical theories, which can then be quantized using standard techniques. This also has implications for the problem of time in quantum gravity and its solution.
@@
The article is divided into five sections and an appendix. In Sec. II, we quickly review the fundamentals of canonical gauge theories (with further details in Appendix A), in order to subsequently present the Fock–Stueckelberg approach to general gauge theories in Sec. III. This section is then subdivided into several particular cases and examples of interest. In Sec. III.2, we discuss how promoting constants of Nature to spacetime fields can be seen as an equivalent approach to the Fock–Stueckelberg theory under certain conditions. The constraint relaxation for “partially Abelianized” systems is then presented in Sec. III.3, and subsequently we turn our attention to mechanical examples in Secs. III.4 and III.5, where we examine the case of time reparametrization invariance and relativistic particles. The Fock–Stueckelberg approach to electromagnetism is presented in Sec. III.6, whereas the application to general relativity is featured in Sec. III.7. Section IV discusses the implications of the Fock–Stueckelberg method to the problem of time in quantum gravity, and we finish with our conclusions in Sec. V.
@@ II Canonical gauge theories in a nutshell
We begin with a brief review of canonical gauge theories before we discuss a general presentation of what we call the Fock–Stueckelberg approach. Their original work on the relativistic particle will be discussed as a particular example of the general framework. We focus on the canonical description of gauge theories because the relation between gauge symmetries and initial-value constraints is most clear in this formulation. As we will see, this leads to a clear implementation of the Fock–Stueckelberg mechanism and its connection to a relaxation of first-class constraints, as well as to the appearance of new degrees of freedom (such as “charge without charge” and “matter without matter”).
@@ II.1 Constraint algebra and gauge transformations
Let us consider a set of fields $z^{m}(\tau,\mathbf{x})$ ($m=1,\ldots,Z$) in $d+1$ spacetime dimensions. We omit the spatial dependence by adopting DeWitt’s compact notation DeWittFields in which abstract indices include not only the discrete values of indices such as $m$, but also the spatial dependence of the fields. For example, $z^{m}(\tau,\mathbf{x})\equiv z^{\mu}(\tau)\equiv z^{\mu}$, [^1] We use the sign $\equiv$ to denote equivalent notations. It is also used when two functions are identically equal (i.e., equal at every point). We will use the sign $\coloneqq$ to denote the definition of a quantity. where $\mu$ is the abstract index that includes both $m$ and $\mathbf{x}$. Repeated indices are then taken to imply a tacit summation over the discrete values of the original indices (e.g., over $m=1,\ldots,Z$) together with a spatial integration. In what follows, as a matter of notational convenience, we also write $\mu=1,\ldots,Z$ and similar expressions for other abstract indices with the understanding that this slight abuse of notation only specifies the discrete values covered by the abstract index. In the case $d=0$ (mechanics), there is of course no spatial dependence and no spatial integrations. We assume that the fields $z^{\mu}$ are bosonic for simplicity. The field-space spanned by $z^{\mu}$ will be referred to as the phase space.
@@
$S[z,\lambda]=\int_{\tau_{0}}^{\tau_{1}}\mathrm{d}\tau\left\{\mathcal{A}_{\mu}[z]\dot{z}^{\mu}-H_{0}(\tau;z]-\lambda^{a}\phi_{a}(\tau;z]\right\}\,,$ (4)
where a dot indicates a derivative with respect to time $\tau$, $\mathcal{A}_{\mu}$ is the symplectic potential, and $\lambda^{a}$ are Lagrange multipliers ($a=1,\ldots,N\leq Z$). The field equations of $z^{\mu}$ are [^3] The derivatives with respect to $z^{\mu}$ are taken in the functional sense if $d>0$. We recall that, for a local functional $\mathcal{F}$ of the field $z(\tau,\mathbf{x})$, which is written as $\mathcal{F}(\tau,\mathbf{x};z]=\int\mathrm{d}^{d}yF\left(\tau,\mathbf{x},\mathbf{y};z(\tau,\mathbf{y}),\partial z(\tau,\mathbf{y}),\ldots,\partial^{n}z(\tau,\mathbf{y})\right)$ with $F$ being a scalar density in $\mathbf{y}$, the functional derivative $\delta\mathcal{F}/\delta z(\tau,\mathbf{x})$ of $\mathcal{F}$ relative to a variation $z(\tau,\mathbf{x})\mapsto z(\tau,\mathbf{x})+\varepsilon\eta(\tau,\mathbf{x})$ is obtained from $\delta\mathcal{F}=\varepsilon\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\mathcal{F}(\tau,\mathbf{x};z+\varepsilon\eta]=:\varepsilon\int\mathrm{d}^{d}y\ \frac{\delta\mathcal{F}}{\delta z(\tau,\mathbf{y})}\eta(\tau,\mathbf{y})$. This leads to $\delta z(\tau,\mathbf{x})=\varepsilon\eta(\tau,\mathbf{x})$, as it should.
@@
are the symplectic form and the total Hamiltonian, respectively. We assume that $\Omega_{\mu\nu}$ is invertible with $\Omega^{\mu\nu}\Omega_{\nu\rho}=\delta^{\mu}_{\rho}$. [^4] If $d>0$, the delta symbol corresponds to a multiplication of a Kronecker delta for the discrete variables and a Dirac delta distribution for the continuous spatial coordinates. Then, Eq. (5) can be rewritten as
@@
In this way, the field equations associated with $z^{\mu}$ can be written compactly as $\dot{z}^{\mu}=\{z^{\mu},H\}$. On the other hand, the field equations associated with the multipliers $\lambda^{a}$ are constraints on the fields:
$\phi_{a}(\tau;z]=0\ .$ (9)
@@
We can now make a series of simplifying assumptions that will be sufficient for our purposes. We assume that $\phi_{a}$ are functions of the fields and their spatial derivatives, $\phi_{a}\equiv\phi_{a}(z,\partial z,\ldots,\partial^{n}z)$ with $n$ finite, and they do not depend explicitly on time,
$\frac{\partial\phi_{a}}{\partial\tau}=0\ .$ (10)
@@
Moreover, the $\phi_{a}$ functions are all functionally independent and not identically zero, so there is no constraint that can be reduced to another (e.g., $\phi_{2}=\phi_{1}^{2}$). The constraint equations (9) together with their first $p$ derivatives ($\partial\phi_{a}=0,\ldots,\partial^{p}\phi_{a}=0$) are also assumed to define a hypersurface $\Sigma$ on the space of the $\mathtt{Z}=(z,\partial z,\ldots,\partial^{n+p}z)$ values, which can be taken to be independent at a fixed arbitrary $\mathbf{x}$. If $\Phi_{\alpha}=0$ are the independent equations among $(\phi_{a}=0,\ldots,\partial^{p}\phi_{a}=0)$, then in open regions covering $\Sigma$, we assume that $\Phi_{\alpha}$ can be regarded as coordinates in a neighborhood of $\Sigma$; i.e., we can define local coordinates $\mathtt{Z}^{\prime}(\mathtt{Z})=(\tilde{\mathtt{Z}}(\mathtt{Z}),\Phi(\mathtt{Z}))$. Then, if $f^{\prime}(\mathtt{Z}^{\prime})\coloneqq f(\mathtt{Z}(\mathtt{Z}^{\prime}))$ is such that $f^{\prime}(\tilde{\mathtt{Z}},\Phi=0)=0$, we obtain [^5] See Theorem 1.1, Appendix 1.A and Theorem 12.4 in HT:book . In Eq. (11), the derivative with respect to $\ell$ is an ordinary partial derivative.
$f^{\prime}(\tilde{\mathtt{Z}},\Phi)$ $=\int_{0}^{1}\mathrm{d}\ell\ \frac{\partial}{\partial\ell}f^{\prime}(\tilde{\mathtt{Z}},\ell\Phi)$ (11)
$=\Phi_{\alpha}\int_{0}^{1}\mathrm{d}\ell\ \frac{1}{\ell}\frac{\partial}{\partial\Phi_{\alpha}}f^{\prime}(\tilde{\mathtt{Z}},\ell\Phi)$
$=:V^{\alpha}(\tilde{\mathtt{Z}},\Phi)\Phi_{\alpha}\ .$
Thus, if a function $f(\mathtt{Z})$ vanishes when the constraints are satisfied, it vanishes on $\Sigma$, and it can be written as a combination of $(\phi_{a},\ldots,\partial^{p}\phi_{a})$ with some finite $p$ and with coefficients that are locally defined by Eq. (11). That this is achievable without the use of boundary conditions (but only with the use of $\Phi_{\alpha}=0$) corresponds to the assumption of ‘local completeness’ of the constraint functions, which, just as the fact that $\Phi_{\alpha}$ serve as local coordinates, constitutes a regularity condition to be imposed on the constraints.
@@
We use Dirac’s ‘weak equality’ sign $\approx$ to denote identities that hold when the constraints are satisfied Dirac2 . [^6] Identities that hold when (all) the field equations are satisfied can be called ‘on-shell identities’. Since every solution to the field equations must satisfy Eqs. (9), then on-shell identities are also weak equalities. In a slight abuse of terminology, we use both weak equalities and on-shell identities as synonyms. In particular, $\phi_{a}\approx 0$. The constraint equations (9) must hold at every instant of time, so that $\dot{\phi}_{a}\approx 0$, otherwise the theory is not consistent. The requirement that $\dot{\phi}_{a}\approx 0$ may lead to new constraints $C_{a}$ that are functionally independent from $\phi_{a}$. In the standard Rosenfeld–Bergmann–Dirac analysis Rosenfeld ; SalisburyRosenfeld ; Dirac3 ; RBD ; HT:book ; Bergmann , the $\phi_{a}$ are called ‘primary’ constraints, as they follow from the structure of the action (4), whereas $C_{a}$ are called ‘secondary’ as they follow from the application of the field equations. [^7] Repeated application of the field equations might lead to ‘tertiary’ constraints and so forth, but all these constraints can be collectively called secondary. For our purposes, it is sufficient to consider that only the primaries $\phi_{a}$ are present; i.e., the conditions $\dot{\phi}_{a}\approx 0$ are satisfied without leading to new (secondary) constraints. In this case, if we demand that $\dot{\phi}_{a}\approx 0$ is satisfied for every arbitrary value of the multipliers $\lambda^{a}$, then $\phi_{a}$ and $H_{0}$ must obey a first-class algebra (one that Poisson-commutes if the constraints hold). Due to Eq. (11), this implies
@@
Is any such transformation allowed? No. If no preferred frame is to be selected, then the field equations must have the same functional form in all frames. This means that $z^{\mu}\to z^{\mu}+\delta z^{\mu}$, $\lambda^{a}\to\lambda^{a}+\delta\lambda^{a}$ should be a symmetry. Then, the Noether charge $G$ must satisfy (see Appendix A):
@@
If $\mathcal{I}$ is gauge invariant, it does not change under arbitrary gauge transformations, i.e., $\delta\mathcal{I}\approx 0$ for any choice of $\varepsilon^{a}$. Using the independence of $\varepsilon^{a}$ and its derivatives implied by the arbitrariness of $\varepsilon^{a}$, we see that Eq. (17) together with $\delta\mathcal{I}\approx 0$ implies the gauge invariance conditions:
@@
A gauge-invariant quantity can be seen as one that is the same in every generalized reference frame. In particular, the constraint functions $\phi_{a}$ are gauge invariant, $\delta\phi_{a}\approx 0$, and thus their functional form does not change under a gauge transformation.